National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 821–840 of 2520Page 42 of 100
NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 821

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Let x be the mean of squares of first n natural numbers and y be the square of mean of first n natural numbers. If \(\frac{\text{x}}{\text{y}} = \frac{55}{42}\), then what is the value of n ?

  1. AOption (a)

    24

  2. BOption (b)

    25

  3. COption (c)

    27

  4. DOption (d)

    30

View answer & explanation

Correct answer

Answer: (c)C.

27

Why this answer is correct

Concept:

  • Mean of x1, x2, ....,xn is \(x_1 + x_2 +...+ x_n \over n\)
  • 1 + 2 + 3 + ....+ n = \(n(n+1) \over 2\)
  • 12 + 22 + 32 + ....+ n2 = \(n(n+1)(2n+1) \over 6\)

Calculation:

Given: x is the mean of squares of first n natural numbers and y is the square of mean of first n natural numbers

⇒ \(x = {1^2 +2^2+3^2+...+ n^2 \over n}\) and  \(y = ({1 +2+3+...+ n \over n})^2\)

⇒  \(x= {{n(n+1)(2n+1) \over 6} \over n}\) and \(y = ({{n(n+1) \over 2} \over n})^2\)

⇒ \(y = {{(n+1)(2n+1) \over 6} }\) and  \(y = {{(n+1)^2 \over 4} }\) 

If \(\frac{\text{x}}{\text{y}} = \frac{55}{42}\), 

⇒ \({ x \over y} = {{ {(n+1)(2n+1) \over 6}\over {(n+1)^2 \over 4} }} = {55 \over 42}\)

⇒ \( { {4(n+1)(2n+1) }\over 6(n+1)^2} = {55 \over 42}\)

⇒ \( { {2(2n+1) }\over 3(n+1)} = {55 \over 42}\)

⇒ 2 × 42(2n + 1) = 3 × 55(n + 1)

⇒ 168n + 84 = 165n + 165

⇒ 3n = 81

⇒ n = 27

∴ The correct option is (3).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 822

National Defence Academy (NDA) 2022• Mathematics• Mathematics

What is the probability of getting a composite number in the list of natural numbers from 1 to 50?

  1. AOption (a)

    \(​\frac{7}{10}\)

  2. BOption (b)

    \(\frac{17}{25}\)

  3. COption (c)

    \(\frac{18}{25}\)

  4. DOption (d)

    \(\frac{33}{50}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{17}{25}\)

Why this answer is correct

Concept:

Probability of an event = \(\text {Favourable outcome } \over \text { Total outcomes}\)

Explanation:

Prime numbers between 1 to 50 are given by,

{2, 3, 5, 7, 11, 13, 17, 19, 23, 29, 31, 37, 41, 43, 47}

⇒ There are 15 prime numbers between 1 to 50

And 1 is neither composite nor prime.

⇒ Number of composite numbers between 1 to 50 = 50 - 15 - 1 = 34

⇒ Favourable outcome = 34

and total outcome = 50

⇒ The probability of getting a composite number in the list of natural numbers from 1 to 50 

= \(\text {Favourable outcome } \over \text { Total outcomes}\)= \(34 \over 50\)

⇒ The probability of getting a composite number in the list of natural numbers from 1 to 50 = \(\frac{17}{25}\)

∴ The correct option is (2).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 823

National Defence Academy (NDA) 2022• Mathematics• Mathematics

If n > 7, then what is the probability that C(n, 7) is a multiple of 7?

  1. AOption (a)

    0

  2. BOption (b)

    \(\frac{1}{7}\)

  3. COption (c)

    \(\frac{1}{2}\)

  4. DOption (d)

    1

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{1}{7}\)

Why this answer is correct

Concept:

Probability of an event = \(\frac{favourable~ number ~of ~outcomes}{total ~no. ~of ~outcomes}\)

C(n, r) = \(\frac{n!}{(n-r)!~r!}\)(where n! = factorial of n)

Calculation:

In the Given question, we need to find the probability that C(n, 7) is a multiple of 7, where n > 7.

So, Finding out favourable number of outcomes i.e.

For n > 7 ,C(n, 7) must be divisible by 7.

Considering n = 8, We get, C(8, 7) = 8 which is not divisible by 7.

Considering n = 9, We get, C(9, 7) = 36 which is not divisible by 7.

Considering n = 10, We get, C(10, 7) = 240 which is not divisible by 7.

Considering n = 11, We get, C(11, 7) = 1980 which is not divisible by 7.

..

..

..

Considering n = 48, We get ,C(48, 7) = which is not divisible by 7.

BUT , On Considering n = 49 , we get C(49, 7) which is divisible by 7 

⇒Before n = 49, C(n, 7) is not divisible by 7.

Thus, favourable number of outcomes = 1.

Total number of outcomes = C(8, 7) + C(9, 7) + ....

Thus, Probability that C(n,7) is divisible by 7 for n>7 = \(\frac{1}{C(8,7)+C(9,7)...}\)= \(\frac{1}{7}\)

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 824

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Two numbers x and y are chosen at random from a set of the first 10 natural numbers. What is the probability that (x + y) is divisible by 4 ?

  1. AOption (a)

    \(\frac{1}{5}\)

  2. BOption (b)

    \(\frac{2}{9}\)

  3. COption (c)

    \(\frac{8}{45}\)

  4. DOption (d)

    \(\frac{7}{45}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{2}{9}\)

Why this answer is correct

Concept:

Probability of an event = \(\frac{favourable~ number ~of ~outcomes}{total ~no. ~of ~outcomes}\)

Calculation:

The set of first ten natural numbers can be visualized as:

A = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10}

Since, We want to find the probability that (x + y) is divisible by 4,

So, for that,

Total number of cases = \(_{}^{10}\textrm{C} \)2 

⇒ Total number of cases = 45

Now,

Number of favourable cases will be the pairs of (x, y) such that (x + y) is divisible by 4. i.e.

Possible pairsDivisible  byTotal number of such pairs
(1, 3)41
(2, 6),(3, 5),(1, 7)83
(2, 10),(3, 9),(5, 7),(4, 8)124
(6, 10),(7, 9)162

⇒Total number of Pairs formed = 1 + 3 + 4 + 2 = 10

Thus,

Probability that (x + y) is divisible by 4 = \(\frac{favourable~ number ~of ~pairs}{total ~no. ~of ~pairs}\)

⇒\(\frac{10}{45}\)

⇒\(\frac{2}{9}\)

Thus, the Probability that (x + y) is divisible by 4 = \(\frac{2}{9}\)

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 825

National Defence Academy (NDA) 2022• Mathematics• Mathematics

A number x is chosen at random from first n natural numbers. What is the probability that the number chosen satisfies x + \(\frac{1}{\text{x}}\) > 2 ?

  1. AOption (a)

    \(\frac{1}{\text{n}}\)

  2. BOption (b)

    \(\frac{1}{(2\text{n})}\)

  3. COption (c)

    \(\frac{(\text{n}−1)}{\text{n}}\)

  4. DOption (d)

    1

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{(\text{n}−1)}{\text{n}}\)

Why this answer is correct

Concept:

Probability of an event = \(\text {Favourable number of outcomes } \over \text {Total number of outcomes}\) 

Calculation:

Given,  x + \(\frac{1}{\text{x}}\) > 2

⇒ x2 - 2x + 1 > 0

⇒ (x - 1)2 > 0

For any x ≠ 1, this inequality is satisfied.

As x is chosen from first n natural numbers.

⇒ Favourable number of choices for x = n - 1

And the total number of choices = n

⇒ Probability for x = \(\frac{(\text{n}−1)}{\text{n}}\)

∴ The correct option is (3).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 826

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Three fair dice are tossed once. What is the probability that they show different numbers that are in AP?

  1. AOption (a)

    \(\frac{1}{12}\)

  2. BOption (b)

    \(\frac{1}{18}\)

  3. COption (c)

    \(\frac{1}{36}\)

  4. DOption (d)

    \(\frac{1}{72}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{1}{18}\)

Why this answer is correct

Concept:

If three numbers are in A.P., then they are of the form a - d, a, a + d, where d is the common difference.

Probability of an event = \(\text {Favourable number of outcomes } \over \text {Total number of outcomes}\) 

Calculation:

Possible numbers in dice = {1, 2, 3, 4, 5, 6}

Total possible in tossing three dice = 6 × 6 × 6 = 216

Case 1: If the common difference AP is 1 or -1

The AP will be (a - 1, a, a + 1)

AP with common difference 1 are (1,2,3), (2,3,4), (3,4,5), (4,5,6), (3,2,1), (4,3,2), (5,4,3), (6,5,4)

So, there can be 4 ways of forming AP with common difference of 4

Case 3: If the common difference AP is 2 or -2

The AP will be (a - 2, a, a + 2)

AP with common difference 2 or -2 are (1,3,5), (2,4,6),(5,3,1), (6,4,2)

So, there can be 2 ways of forming AP with common difference of 2

And the common difference cannot be greater than 2 for forming an AP of 3 numbers.

Favourable outcome (numbers are in AP) = 8 + 4 = 12

⇒ Probability that they show different numbers that are in AP = \(\text {Favourable number of outcomes } \over \text {Total number of outcomes}\)

⇒ Probability that they show different numbers that are in AP = \(12 \over 216\) = \(\frac{1}{18}\)

∴ The correct option is (2).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 827

National Defence Academy (NDA) 2022• Mathematics• Mathematics

If P(A) = 0.5, P(B) = 0.7 and P(A ∩ B) = 0.3, then what is the value of P(A' ∩ B') + P(A' ∩ B) + P(A ∩ B') ?

  1. AOption (a)

    0.6

  2. BOption (b)

    0.7

  3. COption (c)

    0.8

  4. DOption (d)

    0.9

View answer & explanation

Correct answer

Answer: (b)B.

0.7

Why this answer is correct

Concept:

  • P(A') = 1 - P(A)
  • P(A' ∩ B') = P((A ∪ B)') {Probability of not A and not B}
  • P(A and not B) = P(A ∩ B') = P(A) - P(A ∩ B)
  • P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

Calculation:

Given:  P(A) = 0.5, P(B) = 0.7 and P(A ∩ B) = 0.3, 

⇒  P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒  P(A ∪ B) = 0.5 + 0.7 - 0.3 = 0.9__(i)

Now the probability of not A and not B,

P(A' ∩ B') = P((A ∪ B)') = 1 - P(A ∪ B) 

⇒ P(A' ∩ B') = 1 - 0.9 { From (i)}

⇒ P(A' ∩ B') = 0.1 __(ii)

The probability of B and not A,

P(A' ∩ B) = P(B) - P(A ∩ B) 

⇒ P(A' ∩ B) = 0.7 - 0.3 = 0.4 __(iii)

The probability of A and not B,

P(A ∩ B') = P(A) - P(A ∩ B) 

⇒ P(A' ∩ B) = 0.5 - 0.3 = 0.2 __(iv)

From (ii), (iii) and (iv),

P(A' ∩ B') + P(A' ∩ B) + P(A ∩ B') = 0.1 + 0.4 + 0.2 = 0.7

∴ The correct option is (2).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 828

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Five coins are tossed once. What is the probability of getting at most four tails ?

  1. AOption (a)

    \(\frac{31}{32}\)

  2. BOption (b)

    \(\frac{15}{16}\)

  3. COption (c)

    \(\frac{29}{32}\)

  4. DOption (d)

    \(\frac{7}{8}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{31}{32}\)

Why this answer is correct

Concept:

If A and B are independent, then P(A and B) = P(A).P(B)

Explanation:

Probability of getting atmost four tails in five tosses = 1 - Probability of getting all five tails

⇒ Probability of getting atmost four tails in five tosses = 1 - P(T,T,T,T,T)

In one toss, P(H) = \(1 \over 2\) and P(T) = \(1 \over 2\)

Since all five tosses are independent,

⇒ P(T,T,T,T,T) = \(1 \over 2\) × \(1 \over 2\) × \(1 \over 2\) × \(1 \over 2\) × \(1 \over 2\) = \(1 \over 32\)

⇒ Probability of getting atmost four tails in five tosses = 1 - ​​\(1 \over 32\) = ​​​\(31 \over 32\)​

∴ The correct option is (1).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 829

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Three fair dice are thrown. What is the probability of getting a total greater than or equal to 15 ?

  1. AOption (a)

    \(\frac{19}{216}\)

  2. BOption (b)

    \(\frac{1}{12}\)

  3. COption (c)

    \(\frac{17}{216}\)

  4. DOption (d)

    \(\frac{5}{54}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{5}{54}\)

Why this answer is correct

Concept:

Probability of an event = \(\text {Favourable number of outcomes } \over \text {Total number of outcomes}\) 

Calculation:

On tossing a dice, there can be following outcomes = {1,2,3,4,5,6}

If three fair dices are thrown, the sum can be greater than or equal to 15 in the following ways:

(Sum = 15, 16, 17 or 18}

Dice 1Dice 2Dice 3Sum
36615
45615
46515
54615
55515
56415
63615
64515
65415
66315
46616
55616
56516
64616
65516
66416
56617
65617
66517
66618

 

⇒ Number of Favorable outcomes = 20

And the total number of choices when three dices are thrown = 6 × 6 × 6 = 216

⇒ The probability of getting a total greater than or equal to 15 = \(20 \over 216\) = \(\frac{5}{54}\)

∴ The correct option is (4).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 830

National Defence Academy (NDA) 2022• Mathematics• Mathematics

The probability that a person hits a target is 0.5. What is the probability of at least one hit in 4 shots ?

  1. AOption (a)

    \(\frac{1}{8}\)

  2. BOption (b)

    \(\frac{1}{16}\)

  3. COption (c)

    \(\frac{15}{16}\)

  4. DOption (d)

    \(\frac{7}{8}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{15}{16}\)

Why this answer is correct

Concept:

If A and B are independent, then P(A and B) = P(A).P(B)

Explanation:

Given that: Probability that a person hits a target is 0.5

⇒ Probability of not hitting the target = P(not hit) = 1 - 0.5 = 0.5

Now, the probability of at least one hit in 4 shots = 1 - Probability of no hit in four shots

Since the four shots are independent 

⇒ Probability of no hit in four shots = P(not hit) × P(not hit) × P(not hit) × P(not hit) 

⇒ Probability of no hit in four shots = 0.5 × 0.5 × 0.5 × 0,5 = \(( {1 \over 2})^4 = {1 \over 16}\)

⇒ The probability of at least one hit in 4 shots = 1 - \(1 \over 16\) = \(\frac{15}{16}\)

∴ The correct option is (3).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 831

National Defence Academy (NDA) 2022• Mathematics• Mathematics

A box contains 2 white balls, 3 black balls, and 4 red balls. What is the number of ways of drawing 3 balls from the box with at least one black ball?

  1. AOption (a)

    84

  2. BOption (b)

    72

  3. COption (c)

    64

  4. DOption (d)

    48

View answer & explanation

Correct answer

Answer: (c)C.

64

Why this answer is correct

Concept:

Number of ways of choosing r objects out of n objects = \(^nC_r = {n! \over r! (n-r)!}\)

Explanation:

Given that: A box contains 2 white balls, 3 black balls, and 4 red balls

⇒ Box contains total 9 balls

3 balls from the box with at least one black ball can be drawn in the following ways

 

Case 1: 1 black ball 

Number of ways of choosing 1 black ball from 3 black balls and 2 balls from remaining 6 balls = 3C1 × 6C2 = 3 × 15 = 45

Case 2: 2 black balls

Number of ways of choosing 2 black balls from 3 black balls and 1 ball from remaining 6 balls = 3C2 × 6C1 = 3 × 6 = 18

Case 3: 3 black balls

Number of ways of choosing 3 black balls from 3 black balls and 0 ball from remaining 6 balls = 3C3 × 6C0 = 1 × 1 = 1

⇒ Total number of required ways = 45 + 18 + 1 = 64

∴ The correct option is (3).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 832

National Defence Academy (NDA) 2022• Mathematics• Mathematics

During war, one ship out of 5 was sunk on an average in making a certain voyage. What is the probability that exactly 3 out of 5 ships would arrive safely?

  1. AOption (a)

    \(\frac{16}{625}\)

  2. BOption (b)

    \(\frac{32}{625}\)

  3. COption (c)

    \(\frac{64}{625}\)

  4. DOption (d)

    \(\frac{128}{625}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{128}{625}\)

Why this answer is correct

Concept:

Binomial distribution

In n independent trials, if p and q are the probability of success and failure respectively then the probability of exactly r success = nCr pr (q)n-r

Explanation:

Given, During war, one ship out of 5 was sunk on an average in making a certain voyage

⇒ P(sinking) = \(1 \over 5\)  {failure}

⇒ P(safe arriving) =  P(not sinking) = 1 - \(1 \over 5\) = \(4 \over 5\)  {success}

Now the probability that exactly 3 out of 5 ships would arrive safely = Probability of 3 success

 n = 5, p = \(4 \over 5\), q = \(1 \over 5\), r = 3

⇒ The probability that exactly 3 out of 5 ships would arrive safely =  5C3 (\(4 \over 5\))3 (\(1 \over 5\))5-3 

⇒ The probability that exactly 3 out of 5 ships would arrive safely = \({5! \over 3! 2!} ({64\over 125}).({1 \over 25})\) 

⇒ The probability that exactly 3 out of 5 ships would arrive safely = \(\frac{128}{625}\) 

∴ The correct option is (4).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 833

National Defence Academy (NDA) 2022• Mathematics• Mathematics

A card is drawn from a pack of 52 cards. A gambler bets that it is either a spade or an ace. The odds against his winning are

  1. AOption (a)

    9 ∶ 4

  2. BOption (b)

    35 ∶ 17

  3. COption (c)

    17 ∶ 35

  4. DOption (d)

    4 ∶ 9

View answer & explanation

Correct answer

Answer: (a)A.

9 ∶ 4

Why this answer is correct

Concept:

  • Odds against an event = Number of unfavorable choices : Number of favorable choices
  • n(A or B) = n(A ∪ B) = n(A) + n(B) - n(A ∩ B)

Calculation:

Given: A card is drawn from a pack of 52 cards. A gambler bets that it is either a spade or an ace. 

Number of favourable events in winning = n(spade or ace)

⇒ Number of favourable events in winning = n(spade) + n(ace) - n(spade ∩ ace)

∵ in the pack of 52 cards, 1 card is both spade and ace ⇒ n(spade ∩ ace) = 1

⇒ Number of favourable events in winning = 13 + 4 - 1 = 16

⇒ Number of unfavourable events in winning = Total - favourable

⇒ Number of unfavourable events in winning = 52 - 16 = 36

⇒  The odds against winning = Number of unfavorable choices : Number of favorable choices

⇒  The odds against winning = 36 : 16 = 9 : 4

∴ The correct option is (1).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 834

National Defence Academy (NDA) 2022• Mathematics• Mathematics

The coefficient of correlation between ages of husband and wife at the time of marriage for a given set of 100 couples was noted to be 0.7. Assume that all these couples survive to celebrate the silver jubilee of their marriage. The coefficient of correlation at that point of time will be

  1. AOption (a)

    1

  2. BOption (b)

    0.9

  3. COption (c)

    0.7

  4. DOption (d)

    0.3

View answer & explanation

Correct answer

Answer: (c)C.

0.7

Why this answer is correct

Concept:

The correlation coefficient is statistical measure of the strength of the relationship between the relative moments of two variables.

The values range between  -1.0  and  1.0.

It is determined by dividing the covariance by the product of the two variables standard deviation. Thus,

Coefficient of correlation Does not depends on scale and origin.

Calculation:

In the given question, we are given that Coefficient of correlation between the ages of husband and wife at the time of marriage for a set of 100 couples was 0.7

and According to the question, If couples survive to celebrate their silver jubilee, then 

Since, Coefficient of correlation doesn't changes on scale and time

Thus, the Coefficient of correlation at that point will be same as it was earlier 

⇒The Coefficient of correlation = 0.7

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 835

National Defence Academy (NDA) 2022• Mathematics• Mathematics

The completion of a construction job may be delayed due to strike. The probability of strike is 0.6. The probability that the construction job gets completed on time if there is no strike is 0.85 and the probability that the construction job gets completed on time if there is a strike is 0.35. What is the probability that the construction job will not be completed on time ?

  1. AOption (a)

    0.35

  2. BOption (b)

    0.45

  3. COption (c)

    0.55

  4. DOption (d)

    0.65

View answer & explanation

Correct answer

Answer: (b)B.

0.45

Why this answer is correct

Concept:

  • P(A|B) is the probability of A after B happens.
  • P(A ∩ B) + P(A ∩ B̅) = P(A)
  • P(A|B) = \(P(A \cap B)\over P(B)\)
  • P(A̅ ) = 1 - P(A)

Calculation:

Let A be the event that the construction job is completed on time and B be the event that there is strike.

Given:

The probability of strike is 0.6 ⇒ P(B) = 0.6  __(i)

The probability that the construction job gets completed on time if there is no strike is 0.85

⇒ P(A| B̅) = 0.85  __(ii)

The probability that the construction job gets completed on time if there is a strike is 0.35.

⇒ P(A| B) = 0.35 __(iii)

Find: The probability that the construction job will not be completed on time = 1 - P(A)

Solution:

From (ii), 

 P(A| B̅) = 0.85 

⇒ \({P(A \cap \overline B)\over P(\overline B) }= 0.85\)

⇒ \({P(A \cap \overline B) }= 0.85 P(\overline B)\)

⇒ \({P(A \cap \overline B) }= 0.85 (1 -P(B))\) 

⇒ \({P(A \cap \overline B) }= 0.85 (1 - 0.6)\)  {From (i)}

⇒ \({P(A \cap \overline B) }= 0.34\)  __(iv)

Similarly from (iii),

 P(A| B) = 0.35

⇒ \({P(A \cap B)\over P( B) }= 0.35\)

⇒ \({P(A \cap B) }= 0.35 P( B)\)  

⇒ \({P(A \cap B) }= 0.35 (0.6)\)  {From (i)}

⇒ \({P(A \cap B) }= 0.21\)  __(v)

Adding (iv) and (v)

⇒ \({P(A \cap \overline B) }+ {P(A \cap B) }= 0.34 + 0.21\)

⇒ P(A) = 0.55

⇒ The probability that the construction job is not completed on time = 1 - P(A) = 1 - 0.55

⇒ The probability that the construction job is not completed on time = 0.45

∴ The correct option is (2).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 836

National Defence Academy (NDA) 2022• Mathematics• Mathematics

The mean and standard deviation (SD) of marks obtained by 50 students of a class in 4 subjects are given below :

SubjectMathematicsPhysicsChemistryBiology
Mean Marks40283836
SD15121416

Which one of the following subjects shows highest variability of marks ?

  1. AOption (a)

    Mathematics

  2. BOption (b)

    Physics

  3. COption (c)

    Chemistry

  4. DOption (d)

    Biology

View answer & explanation

Correct answer

Answer: (d)D.

Biology

Why this answer is correct

Concept:

The data with the greatest coefficient of variation has the highest variability.

Coefficient of variation = \({S.D. \over \text {Mean}} \times 100\)

where S.D. is the standard deviation of the data.

Calculation:

Given:

SubjectMathematicsPhysicsChemistryBiology
Mean Marks40283836
SD15121416

 

So, the coefficient of variation for Mathematics = \({15 \over \text 40} \times 100 = {75 \over 2} \)

⇒ The coefficient of variation for Mathematics = 37.5  __(i)

The coefficient of variation for Physics = \({12 \over \text 28} \times 100 = 42.85\) __(ii)

 The coefficient of variation for Chemistry = \({14 \over \text 38} \times 100 = 36.84\)____(iii)

The coefficient of variation for Biology = \({16 \over \text 36} \times 100 = 44.44\)____(iv)

Comparing (i), (ii), (iii) and (iv),

The subject Biology has the greatest Coefficient of Variation

⇒ Biology shows the highest variability.

The correct option is (4).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 837

National Defence Academy (NDA) 2022• Mathematics• Mathematics

The mean and standard deviation (SD) of marks obtained by 50 students of a class in 4 subjects are given below :

SubjectMathematicsPhysicsChemistryBiology
Mean Marks40283836
SD15121416

What is the coefficient of variation of marks in Mathematics ?

  1. AOption (a)

    37.5%

  2. BOption (b)

    38.0%

  3. COption (c)

    38.5%

  4. DOption (d)

    39.0%

View answer & explanation

Correct answer

Answer: (a)A.

37.5%

Why this answer is correct

Concept:

The data with the greatest coefficient of variation has the highest variability.

Coefficient of variation = \({S.D. \over \text {Mean}} \times 100\)

where S.D. is the standard deviation of the data.

Calculation:

Given:

SubjectMathematicsPhysicsChemistryBiology
Mean Marks40283836
SD15121416

 

⇒ S.D. of mathematics = 15 and Mean for Mathematics = 40

So, the coefficient of variation for Mathematics =  \({15 \over \text 40} \times 100 = {75 \over 2} \)

⇒ The coefficient of variation for Mathematics = 37.5  

∴ The correct option is (1).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 838

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Consider the following grouped frequency distribution :

Class0-1010-2020-3030-4040-5050-60
Frequency124643

What is the median of the distribution ?

  1. AOption (a)

    34

  2. BOption (b)

    34.5

  3. COption (c)

    35

  4. DOption (d)

    35.5

View answer & explanation

Correct answer

Answer: (c)C.

35

Why this answer is correct

Concept:

For a distribution, Median = \(L + {{N \over 2} - CF \over f} \times h\)

where, L = lower class limit of median class

N = Sum of frequencies

CF = cumulative frequency of class preceding the median class

f = frequency of median class

h = class length of median class.

Calculation:

ClassFrequencyCumulative frequency
0-1011
10-2023
20-3047
30-40613
40-50417
50-60320
∑ f = N = 20

 

∑f = N = 20

⇒ N/2 = 10 

As 7 < 10 < 13, N/2 lies in the class length 30 - 40 by comparing cumulative frequency.

Median class = 30 - 40

⇒ L = 30, N = 60, CF = 7, f = 6, h = 40 - 30 - 10

⇒ Median = \(L + {{N \over 2} - CF \over f} \times h = 30 + {{20 \over 2} - 7 \over 6} \times 10\)

⇒ Median = \(30+ {3 \over 6} \times 10\)= 35

∴ The correct option is (3).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 839

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Consider the following grouped frequency distribution :

Class0-1010-2020-3030-4040-5050-60
Frequency124643

What is mean deviation about the median ?

  1. AOption (a)

    11.4

  2. BOption (b)

    11.1

  3. COption (c)

    10.8

  4. DOption (d)

    10.5

View answer & explanation

Correct answer

Answer: (d)D.

10.5

Why this answer is correct

Concept:

For a distribution, Median =  \(L + {{N \over 2} - CF \over f} \times h\)

where, L = lower class limit of median class

N = Sum of frequencies

CF = cumulative frequency of class preceding the median class

f = frequency of median class

h = class length of median class.

Mean deviation about median = \({∑ f_i|x_i -M | \over N}\)

where M is median, N is sum of frequencies, xi are the mid points of the class and fi are the frequencies

Calculation:

ClassFrequencyCumulative frequency
0-1011
10-2023
20-3047
30-40613
40-50417
50-60320
∑ f = N = 20

 

∑f = N = 20

⇒ N/2 = 10 

As 7 < 10 < 13, N/2 lies in the class length 30 - 40 by comparing cumulative frequency.

Median class = 30 - 40

⇒ L = 30, N = 60, CF = 7, f = 6, h = 40 - 30 - 10

⇒ Median = \(L + {{N \over 2} - CF \over f} \times h = 30 + {{20 \over 2} - 7 \over 6} \times 10\)

⇒ Median = \(30+ {3 \over 6} \times 10\)= 35

Now calculating mean deviation about median,

ClassFrequency(fi)Midpoint of class (xi)(xi - M)fi|xi - M|
0-10155 - 35 = -3030
10-2021515 - 35 = -2040
20-3042525 - 35 = -1040
30-4063535 - 35 = 00
40-5044545 - 35 = 1040
50-6035555 - 35 = 2060
∑f = N = 20∑fi(xi - M) = 210

 

⇒ Mean deviation about median = \({∑ f_i|x_i -M | \over N}\) = \(210\over 20 \) = 10.5

∴ The correct option is (4).

NDA Mathematics 4 Sep 2022 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 840

National Defence Academy (NDA) 2022• Mathematics• Mathematics

Consider the following grouped frequency distribution :

Class0-1010-2020-3030-4040-5050-60
Frequency124643

What is the mean deviation about the mean ?

  1. AOption (a)

    10.15

  2. BOption (b)

    10.65

  3. COption (c)

    11.15

  4. DOption (d)

    11.65

View answer & explanation

Correct answer

Answer: (b)B.

10.65

Why this answer is correct

Concept:

  • Mean of a distribution = \(∑ f_i x_i \over ∑ f_i\)
  • Mean deviation about mean x̅ = \({∑ f_i|x_i -\overline x | \over ∑ f_i}\)

where fi are frequencies and xi are the mid points of class

Calculation:

Calculating mean

ClassFrequency(fi)Midpoint of class (xi)fi xi
0-10155
10-2021530
20-30425100
30-40635210
40-50445180
50-60355165
∑fi = N = 20∑fixi = 690

 

⇒ Mean = \(∑ f_i x_i \over ∑ f_i\) = \(690 \over 20\) = 34.5

 

Calculating the mean deviation about mean

ClassFrequency(fi)Midpoint of class (xi)|xi - x̅ |fi|xi - x̅|
0-1015|5 - 34.5| = 29.529.5
10-20215|15 - 34.5| = 19.539
20-30425|25 - 34.5| = 9.538
30-40635|35 - 34.5| = 0.53
40-50445|45 - 34.5| = 10.542
50-60355|55 - 34.5| = 20.561.5
∑fi = N = 20∑fi|xi - x̅| = 213

 

⇒ Mean deviation about mean x̅ = \({∑ f_i|x_i -\overline x | \over ∑ f_i}\) = \(213 \over 20\) = 10.65

∴ The correct option is (2).

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.